OpenStudy (studytheworld):

Help! Will medal and fan! please helpp Suppose a triangle has sides a, b and c, and that a^2+b^2

OpenStudy (studytheworld):

1. (theta) is an obtuse triangle. 2. The triangle is not a right triangle. 3. The triangle is a right triangle. 4. cos(theta) < 0

OpenStudy (studytheworld):

@TrojanPoem Could you help on this one, too?

OpenStudy (trojanpoem):

 \[\cos \theta = \frac{ a^2 + b^2 - c^2 }{ 2ab }\] a^2 + b^2 - c^2 < 0 costheta * 2ab = a^2 + b^2 - c^2 2abcostheta < 0 ( b, a are postive so we won't flip the sign) costheta < 0

OpenStudy (studytheworld):

Okay I see. So answer 4 for sure

OpenStudy (studytheworld):

Answer 2, as well?

OpenStudy (trojanpoem):

the triangle is obtuse too, as a^2 + b^2 < c^2 Notice that : a2 + b2 > c2 (acute) a2 + b2 = c2 ( right) a2 + b2 < c2 ( obtuse)

OpenStudy (studytheworld):

Okay, awesome! Thanks so much for helping me again!

OpenStudy (trojanpoem):

Any time.

OpenStudy (trojanpoem):

The triangle isn't a right triangle as it's obtuse.